What is the Laplace transform of t^3 + cos3t-e^-3t?

1
Expert's answer
2022-02-07T16:05:39-0500

We have "F_{(t)}= {t^3 + \cos{(3t)}-e^{-3t}}"


We proceed to use the rules for Laplace transform and we proceed with each one of the terms:


"\mathscr{L}[F_{(t)}]_{(s)}=\mathscr{L}[{t^3 + \cos{(3t)}-e^{-3t}}]_{(s)}
\\ \mathscr{L}[F_{(t)}]_{(s)}=\mathscr{L}[{t^3}]_{(s)}+\mathscr{s}[{\cos{(3t)}}]_{(s)}-\mathscr{L}[{e^{-3t}}]_{(s)}
\\ \mathscr{L}[F_{(t)}]_{(s)}=\dfrac{3!}{s^{3+1}}+\dfrac{s}{s^2+3^2}-\dfrac{1}{s-{(-3)}}
\\ \therefore \mathscr{L}[F_{(t)}]_{(s)}=\dfrac{6}{s^{4}}+\dfrac{s}{s^2+9}-\dfrac{1}{s+3}"


In conclusion, the Laplace transform is "\mathscr{L}[F_{(t)}]_{(s)}=\dfrac{6}{s^{4}}+\dfrac{s}{s^2+9}-\dfrac{1}{s+3}".